# UAV-Mounted IRS (UMI) in the Presence of Hovering Fluctuations: 3D Pattern Characterization and Performance Analysis

### v2

@@ III IRS Radiation pattern Distribution Function
Utilizing distribution of $\mathbb{G}_{a}$ and $\mathbb{G}_{e}$ in (30) and (36), respectively, the distribution of $\mathbb{G}$ can be obtained.
@@
where $J=J_{a}\times J_{e}$, $\mathbf{q}=\mathbf{q}_{e}\otimes\mathbf{q}_{a}$ and $\mathbf{p}=\mathbf{p}_{e}\otimes\mathbf{p}_{a}$.
@@
As $\mathbb{G}$ is equal to the product of $\mathbb{G}_{e}$ and $\mathbb{G}_{a}$, the PDF of $\mathbb{G}$ can be obtained using the same steps outlined in the proof of Lemma 4 in Appendix C. ∎
@@ IV Performance Analysis
This section conducts a comprehensive performance analysis for four scenarios: SISO-Passive-IRS, MISO-Passive-IRS, SISO-Active-IRS, and MISO-Active-IRS. The closed-form expressions for each scenario are presented when taking into account the fluctuation of the IRS mounted on UAV. Since the rate of both UAV fluctuations and channel variations is less than the symbol rate, we will focus on the OP metric that is specified as
@@
IV-A SISO-Passive-IRS
The SNR of SISO-Passive-IRS system can be calculated by rewriting (12) as
$\gamma=\frac{P_{t}\mathbb{G}|\bar{\mathbf{h}}^{H}_{IU}\mathbf{\Phi}\bar{\mathbf{h}}_{BI}|^{2}}{N_{0}},$ (41)
where $\bar{\mathbf{h}}_{BI}$ is $\bar{\mathbf{H}}_{BI}$ when $M=1$. It is known that the maximum SNR is achieved when the phase shift is adjusted as $\phi_{n}=\angle[\bar{\mathbf{h}}_{IU}]_{n}-\angle[\bar{\mathbf{h}}_{BI}]_{n},~{}\forall n$. Hence, the maximum instantaneous SNR is evaluated as
$\gamma_{\text{max}}=\mathbb{G}\gamma_{0}(\sum_{n=1}^{N}{|\bar{h}_{IU,n}||\bar{h}_{BI,n}|})^{2},$ (42)
where $\gamma_{0}=\frac{P_{t}}{N_{0}}$. We define $U=V^{2}$ where
$V=\sum_{n=1}^{N}{|\bar{h}_{IU,n}||\bar{h}_{BI,n}|}.$ (43)
From (40) and (42), the OP can be obtained as
$P_{\text{out}}$ $=\text{Pr}(\mathbb{G}U\leq\frac{\gamma_{th}}{\gamma_{0}})$ (44)
[…]
where $F_{U}(.)$ is the CDF of random variable $U$. The PDF of $\mathbb{G}$, $f_{\mathbb{G}}(\mathbb{G})$ is derived in Theorem 1, i.e. in (39). By substituting (39) into (44), the OP can be given as
$P_{\text{out}}=\sum_{i=1}^{J}p(i)F_{U}(\frac{\gamma_{th}}{\gamma_{0}q(i)}),$ (45)
To determine $F_{U}(.)$ in (45), we estimate $V$, given in (43) which is the mixture channels associated with IRS by using the CLT and Gamma approximation.
@@ IV-A SISO-Passive-IRS
Under CLT approximation, the closed-form OP for the SISO-Passive-IRS system is given by
$P_{\text{out}}=\sum_{i=1}^{J}p(i)Q(\frac{\mu_{v}-\sqrt{\frac{\gamma_{th}}{\gamma_{0}q(i)}}}{\sigma_{v}}),$ (46)
@@
$\mu_{v}=\frac{N\pi c_{0}}{4\sqrt{(K_{0}+1)(K_{1}+1)}d_{0}^{\alpha_{0}/2}d_{1}^{\alpha_{1}/2}}L_{\frac{1}{2}}(-K_{0})L_{\frac{1}{2}}(-K_{1}),$ (47)
$\sigma^{2}_{v}=\frac{c_{0}^{2}N}{d_{0}^{\alpha_{0}}d_{1}^{\alpha_{1}}}$ $-\frac{\pi^{2}c_{0}^{2}N}{16(K_{0}+1)(K_{1}+1)d_{0}^{\alpha_{0}}d_{1}^{\alpha_{1}}}$ (48)
×$\quad\times(L_{\frac{1}{2}(-K_{0})})^{2}\times(L_{\frac{1}{2}(-K_{1})})^{2}.$
@@
Since the value of $Q(\frac{\mu_{v}+\sqrt{u}}{\sigma_{v}})$ is negligible rather than $Q(\frac{\mu_{v}-\sqrt{u}}{\sigma_{v}})$, the CDF of $U$ is approximated as
$F_{U}(u)\approx Q(\frac{\mu_{v}-\sqrt{u}}{\sigma_{v}}).$ (52)
Finally, by substituting (52) into (45), $P_{out}$ can be obtained. ∎
Considering the Gamma distribution for $V$, we can derive the closed-form OP for the SISO-Passive-IRS system given in Theorem 3.
@@
In the case of Gamma approximation, the OP of the SISO-Passive-IRS system can be provided as follows
$P_{\text{out}}=\sum_{i=1}^{J}p(i)(\frac{\gamma(\Lambda,\sqrt{\frac{\gamma_{th}}{\gamma_{0}q(i)}}/\Omega)}{\Gamma(\Lambda)}),$ (53)
@@
where $\Lambda=\frac{\mathbb{E}^{2}(V)}{\text{var}(V)}$ and $\Omega=\frac{\text{var}(V)}{\mathbb{E}(V)}$. By substituting (54) into (45), $P_{out}$ is derived. ∎
IV-B MISO-Passive-IRS
According to the discussion in [32] and [33], we can consider only the LOS part in the MISO system due to the strong directivity of multiple antennas at the BS and the negligible NLOS component. Hence, the maximum instantaneous SNR is evaluated as
$\gamma_{\text{max}}=M\gamma_{0}\mathbb{G}(\sqrt{\beta_{0}}\sum_{n=1}^{N}{|\bar{h}_{IU,n}|})^{2}$ (55)
Suppose $V_{2}=\sqrt{\beta_{0}}\sum_{n=1}^{N}{|\bar{h}_{IU,n}|}$ and $U_{2}=V_{2}^{2}$. The OP can be obtained as
$P_{\text{out}}$ $=\text{Pr}(\mathbb{G}U_{2}\leq\frac{\gamma_{th}}{M\gamma_{0}})$ (56)
[…]
By substituting (39) into (56), the OP is derived as
$P_{\text{out}}=\sum_{i=1}^{J}p(i)F_{U_{2}}(\frac{\gamma_{th}}{M\gamma_{0}q(i)}),$ (57)
where $F_{U_{2}}(.)$ is the CDF of random variable $U_{2}$. The PDF of $\mathbb{G}$, $f_{\mathbb{G}}(\mathbb{G})$ derived in Theorem 1 in (39). In this section similar to the previous section, we approximate the distribution of $V_{2}$ using CLT and Gamma approximations.
Proposition 1.
According to CLT approximation, the closed-form OP for the MISO-Passive-IRS system is obtained as
$P_{\text{out}}=\sum_{i=1}^{J}p(i)Q(\frac{\mu_{v_{2}}-\sqrt{\frac{\gamma_{th}}{M\gamma_{0}q(i)}}}{\sigma_{v_{2}}}),$ (58)
where
$\mu_{v_{2}}=\frac{N\sqrt{\pi}c_{0}}{2\sqrt{(K_{1}+1)}d_{0}^{\alpha_{0}/2}d_{1}^{\alpha_{1}/2}}L_{\frac{1}{2}}(-K_{1})$ (59)
$\sigma^{2}_{v_{2}}=\frac{c_{0}^{2}N}{d_{0}^{\alpha_{0}}d_{1}^{\alpha_{1}}}-\frac{N\pi c_{0}^{2}}{4(K_{1}+1)d_{0}^{\alpha_{0}}d_{1}^{\alpha_{1}}}\times(L_{\frac{1}{2}(-K_{1})})^{2}$ (60)
Proof.
It can be proven similar to Theorem 2, by replacing variable $V$ and $\gamma_{0}$ with $V_{2}$ and $M\gamma_{0}$, respectively. Additionally, the mean and variance of RV $V_{2}$ can be calculated using basic mathematical operations. ∎
We can also consider the gamma distribution for $V_{2}$, and therefore calculate the closed-form OP for the MISO-Passive-IRS system as given in Proposition 2.
Proposition 2.
In the case of Gamma approximation, the OP of the MISO-Passive-IRS system is obtained as follows
$P_{\text{out}}=\sum_{i=1}^{J}p(i)(\frac{\gamma(\Lambda_{2},\sqrt{\frac{\gamma_{th}}{M\gamma_{0}q(i)}}/\Omega_{2})}{\Gamma(\Lambda_{2})}),$ (61)
where $\Lambda=\frac{\mu^{2}_{V_{2}}}{\sigma^{2}_{V_{2}}}$ and $\Omega=\frac{\sigma_{V_{2}}^{2}}{\mu_{V_{2}}}$. $\mu_{V_{2}}$ and $\sigma^{2}_{V_{2}}$ are given by (59) and (60), respectively.
Proof.
It can be proven similar to Theorem 3, by replacing variable $V$, $\Lambda$, $\Omega$ and $\gamma_{0}$ with $V_{2}$, $\Lambda_{2}$, $\Omega_{2}$ and $M\gamma_{0}$, respectively. ∎
IV-C SISO-Active-IRS
Based on the discussion in [31], for simplicity, we assume that each reflecting elements has the same amplification factor $A$. Consequently, the optimal active-IRS reflection design can be derived as follows
$[\mathbf{\Phi}^{*}]_{n}=e^{j(\angle[\bar{\mathbf{h}}_{IU}]_{n}-\angle[\bar{\mathbf{h}}_{BI}]_{n})},~{}\forall n,$ (62)
$(A^{*})^{2}=\frac{P_{F}}{P_{t}|\bar{\mathbf{h}}_{BI}|^{2}+NN_{F}}.$ (63)
Substituting (62) and (63) into (12), the SNR of SISO-Active-IRS system can be expressed as
$\gamma_{\text{max}}=\frac{(\sum_{n=1}^{N}{|\bar{h}_{IU,n}||\bar{h}_{BI,n}|})^{2}}{c_{1}\underbrace{\sum_{n=1}^{N}|\bar{h}_{IU,n}|^{2}}_{Z_{0}}+c_{2}/\mathbb{G}\underbrace{\sum_{n=1}^{N}|\bar{h}_{BI,n}|^{2}}_{Z_{1}}+c_{3}/\mathbb{G}},$ (64)
where $c_{1}=\frac{N_{F}}{P_{t}}$, $c_{2}=\frac{N_{0}}{P_{F}}$ and $c_{3}=\frac{NN_{0}N_{F}}{P_{t}P_{F}}$. Under CLT approximation, $Z_{k},~{}k\in\{0,1\}$, defined in (64), follows Gaussian distribution, specifically $Z_{k}\sim\mathcal{N}(\mu_{Z_{k}},\sigma^{2}_{Z_{k}})$ with mean and variance as follows
$\mu_{Z_{k}}$ $=\frac{Nc_{0}}{d_{k}^{\alpha_{k}}},$ (65)
$\sigma^{2}_{Z_{k}}$ $=\frac{N(K_{k}^{2}+4K_{k}+2)c_{0}^{2}}{(K_{k}+1)^{2}d_{k}^{2\alpha_{k}}}-\frac{Nc_{0}^{2}}{d_{k}^{2\alpha_{k}}},~{}\forall k$
@@ IV-C SISO-Active-IRS
In the case of CLT approximation, the OP of the SISO-Active-IRS system can be obtained as
@@
$R(\mathbb{G})=Q(\frac{\mu_{v}-\sqrt{\mu_{w}(\mathbb{G})}}{\sigma_{v}})+F_{U}^{(2)}(\mu_{w}(\mathbb{G}))\sigma_{w}^{2}(\mathbb{G}),$ (67)
$F_{U}^{(2)}(\mu_{w}(\mathbb{G}))$ $=\frac{1}{{4\sqrt{2\pi}\sigma_{v}^{3}}}\exp(\frac{-(\mu_{v}-\sqrt{\mu_{w}(\mathbb{G})})^{2}}{2\sigma^{2}_{v}})$ (68)
×$\times({\frac{\mu_{v}}{\mu_{w}(\mathbb{G})}-\frac{1}{(\mu_{w}(\mathbb{G}))^{\frac{1}{2}}}-\frac{\sigma_{v}^{2}}{(\mu_{w}(\mathbb{G}))^{\frac{3}{2}}}}).$
Here, $\mu_{w}(\mathbb{G})=c_{1}\gamma_{th}\mu_{Z_{0}}+\frac{c_{2}\gamma_{th}}{\mathbb{G}}\mu_{Z_{1}}+\frac{c_{3}\gamma_{th}}{\mathbb{G}}$ and $\sigma^{2}_{w}(\mathbb{G})=\gamma_{th}^{2}c^{2}_{1}\sigma^{2}_{Z_{0}}+\frac{c_{2}^{2}\gamma_{th}^{2}}{\mathbb{G}^{2}}\sigma^{2}_{Z_{1}}$
@@
Let $Z(\mathbb{G})=c_{1}Z_{0}+c_{2}Z_{1}/\mathbb{G}$. Since two RVs $Z_{0}$ and $Z_{1}$ are independent, the RV of $Z(\mathbb{G})$ follows Gaussian distribution when RV of $\mathbb{G}$ is known i.e. $Z(\mathbb{G})\sim\mathcal{N}(\mu_{Z}(\mathbb{G}),\sigma^{2}_{Z}(\mathbb{G}))$, which $\mu_{Z}(\mathbb{G})=c_{1}\mu_{Z_{0}}+c_{2}/\mathbb{G}\mu_{Z_{1}}$ and $\sigma^{2}_{Z}(\mathbb{G})=c_{1}^{2}\sigma^{2}_{Z_{0}}+c^{2}_{2}\mathbb{G}^{2}\sigma^{2}_{Z_{1}}$. Now, we can rewrite SNR in (64) as
$\gamma_{max}=\frac{U}{Z(\mathbb{G})+c_{3}/\mathbb{G}},$ (69)
where $U=\sum_{n=1}^{N}{|\bar{h}_{IU,n}||\bar{h}_{BI,n}|}$. Therefore, the OP can be given as
[…]
@@
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})$ $=\text{Pr}(\frac{U}{Z(\mathbb{G})+c_{3}/\mathbb{G}}\leq\gamma_{\text{th}})$ (71)
[…]
$=\mathbb{E}_{W}\{F_{U}(w)\},$
where $W(\mathbb{G})=\gamma_{th}(Z+c_{3}/\mathbb{G})$. The distribution of $W$ follows $\mathcal{N}(\mu_{w}(\mathbb{G}),\sigma_{w}^{2}(\mathbb{G}))$ where $\mu_{w}(\mathbb{G})=c_{1}\gamma_{th}\mu_{Z_{0}}+\frac{c_{2}\gamma_{th}}{\mathbb{G}}\mu_{Z_{1}}+\frac{c_{3}\gamma_{th}}{\mathbb{G}}$ and $\sigma^{2}_{w}(\mathbb{G})=\gamma_{th}^{2}c^{2}_{1}\sigma^{2}_{Z_{0}}+\frac{c_{2}^{2}\gamma_{th}^{2}}{\mathbb{G}^{2}}\sigma^{2}_{Z_{1}}$ From (52), we have
$F_{U}(w(\mathbb{G}))=Q(\frac{\mu_{v}-\sqrt{w(\mathbb{G})}}{\sigma_{v}}).$ (72)
For a function of one variables $f(x)$ whose first- and second-order derivatives $f^{(1)}(x)$ and $f^{(2)}(x)$, respectively, exist at the point $a$, the $2^{\footnotesize\text{st}}$-order Taylor polynomial near the point $a$ is
$f(x)\approx f(a)+f^{(1)}(a)(x-a)+f^{(2)}(a)(x-a)^{2}.$ (73)
Using (73), we approximate $F_{U}(w)$ in (72) at point $\mu_{w}(\mathbb{G})$ as follows
$F_{U}(w(\mathbb{G}))$ $\approx F_{U}(\mu_{w}(\mathbb{G}))$ (74)
$+F_{U}^{(1)}(\mu_{w}(\mathbb{G}))(w(\mathbb{G})-\mu_{w}(\mathbb{G}))$
$+F_{U}^{(2)}(\mu_{w}(\mathbb{G}))(w(\mathbb{G})-\mu_{w}(\mathbb{G}))^{2}.$

### v3

@@ III IRS Radiation Pattern Distribution Function
Utilizing distribution of $\mathbb{G}_{a}$ in (37) and treated expression of $\mathbb{G}_{e}$ in (44), the distribution of $\mathbb{G}$ can be obtained.
@@
where $J=l^{2}D^{2}$, $\mathbf{q}=q_{e}\mathbf{q}_{a}$, $q_{e}=\cos^{3}(\theta_{t})\times\cos^{3}(\theta_{r})$, and $\mathbf{p}=\mathbf{p}_{a}$.
@@
As $\mathbb{G}$ is equal to the product of $\mathbb{G}_{e}$ and $\mathbb{G}_{a}$, the PDF of $\mathbb{G}$ is calculated as $f_{\mathbb{G}}(\mathbb{G})=\frac{1}{\mathbb{G}_{e}}f_{\mathbb{G}_{a}}(\mathbb{G}/\mathbb{G}_{e})$. By substituting (37) and (44), the PDF of $\mathbb{G}_{e}$ is obtained. ∎
@@ IV Performance Analysis
This section conducts a comprehensive performance analysis for passive/active IRS. The closed-form expressions for both scenarios are presented when taking into account the fluctuation of the IRS mounted on the UAV. Since the rate of both UAV fluctuations and channel variations is less than the symbol rate, we will focus on the OP metric that is specified as
@@
IV-A Passive-IRS
The instantaneous maximum SNR of the Passive-IRS system can be calculated by rewriting (19) as
$\gamma_{\text{max}}=\frac{M\gamma_{0}\mathbb{G}U}{b_{1}},$ (47)
where $\gamma_{0}=\frac{P_{t}}{\sigma_{n}^{2}}$ and $b_{1}=\frac{P_{t}\zeta\sigma_{e}^{2}}{\sigma_{n}^{2}}+1$. Here, we define $U=V^{2}$ where $V=\sqrt{\beta_{0}\beta_{1}(1-\zeta)}\sum_{n=1}^{N}|H_{BI,m,n}||h_{IU,n}|$. From (46) and (47), the OP can be obtained as
$P_{\text{out}}$ $=\text{Pr}(\mathbb{G}U\leq\frac{b_{1}\gamma_{th}}{M\gamma_{0}})$ (48a)
$=\int\text{Pr}(\mathbb{G}U\leq\frac{b_{1}\gamma_{th}}{M\gamma_{0}}|\mathbb{G})f_{\mathbb{G}}(\mathbb{G})d\mathbb{G}$ (48b)
$=\int F_{U}(\frac{b_{1}\gamma_{th}}{M\gamma_{0}\mathbb{G}})f_{\mathbb{G}}(\mathbb{G})d\mathbb{G},$ (48c)
where $F_{U}(.)$ is the CDF of random variable $U$. The PDF of $\mathbb{G}$, $f_{\mathbb{G}}(\mathbb{G})$ is derived in Theorem 1 (Eq. 45). By substituting (45) into (48), the OP can be given as
$P_{\text{out}}=\sum_{i=1}^{J}p(i)F_{U}(\frac{b_{1}\gamma_{th}}{M\gamma_{0}q(i)}),$ (49)
To determine the CDF of $U$, $F_{U}(.)$ in (49), we estimate $V$, which is the mixture channels associated with IRS, by using the CLT and Gamma approximation.
@@ IV-A Passive-IRS
Under CLT approximation, the closed-form OP for the Passive-IRS system is given by
$P_{\text{out}}=\sum_{i=1}^{J}p(i)T(q(i)),$ (50)
@@
$T(x)=Q(\frac{\mu_{v}-\sqrt{\frac{b_{1}\gamma_{th}}{M\gamma_{0}x}}}{\sigma_{v}})-Q(\frac{\mu_{v}+\sqrt{\frac{b_{1}\gamma_{th}}{M\gamma_{0}x}}}{\sigma_{v}}),$ (51)
$\mu_{v}=\frac{N\pi c_{0}\sqrt{1-\zeta}}{4\sqrt{(K_{0}+1)(K_{1}+1)}d_{0}^{\alpha_{0}/2}d_{1}^{\alpha_{1}/2}}L_{\frac{1}{2}}(-K_{0})L_{\frac{1}{2}}(-K_{1}),$ (52)
$\sigma^{2}_{v}=\frac{c_{0}^{2}N(1-\zeta)}{d_{0}^{\alpha_{0}}d_{1}^{\alpha_{1}}}(1$ $-\frac{\pi^{2}}{16(K_{0}+1)(K_{1}+1)}$ (53)
$\quad\quad\quad\times(L_{\frac{1}{2}(-K_{0})})^{2}\times(L_{\frac{1}{2}(-K_{1})})^{2}).$
@@
Finally, by substituting (56) into (49), the OP can be obtained. ∎
Considering the Gamma distribution for $V$, we can derive the closed-form OP for the Passive-IRS given in Theorem 3.
@@
In the case of Gamma approximation, the OP of the Passive-IRS system can be provided as follows
$P_{\text{out}}=\sum_{i=1}^{J}p(i)(\frac{\gamma(\Lambda,\sqrt{\frac{b_{1}\gamma_{th}}{M\gamma_{0}q(i)}}/\Omega)}{\Gamma(\Lambda)}),$ (57)
@@
where $\Lambda=\frac{\mathbb{E}^{2}(V)}{\text{var}(V)}$ and $\Omega=\frac{\text{var}(V)}{\mathbb{E}(V)}$. By substituting (58) into (49), the OP is derived. ∎
Remark 3.
For a large number of passive IRS elements ($N\rightarrow\infty$), the scaling $\mathbb{G}\propto 1/N^{2}$ is counteracted by $U\propto N^{2}$. Consequently, these scaling laws cancel, rendering the maximum SNR, $\gamma_{\text{max}}=\frac{M\gamma_{0}\mathbb{G}U}{b_{1}}$, independent of $N$ under perfect CSI. As a result, the maximum SNR $\gamma_{\text{max}}$ saturates at a fixed value, independent of $N$. The OP thus converges to a constant floor dictated by UAV fluctuations, resulting in a diversity order of zero. This is in contrast to the case without fluctuations, which achieves an infinite diversity order under perfect CSI. Under imperfect CSI, the OP for both scenarios (with and without fluctuations) converges to a constant floor dictated by the CSI error factor.
IV-B Active-IRS
The instantaneous maximum SNR of the Active-IRS system can be obtained by rewriting (15) as
$\gamma_{\text{max}}=\frac{M\gamma_{0}\mathbb{G}U}{c_{1}Z_{0}+c_{2}Z_{1}+c_{3}},$ (59)
where $\gamma_{0}$ and $U$ defined in (47). Here, $c_{1}=\frac{\sigma_{f}^{2}}{\sigma_{n}^{2}}\cos^{3}(\theta_{r})$, $c_{2}=\frac{P_{t}}{P_{F}}(1+\frac{P_{t}\zeta\sigma_{e}^{2}}{\sigma_{n}^{2}})\cos^{3}(\theta_{t})$, $c_{3}=\frac{N\sigma_{f}^{2}}{P_{F}}(1+\frac{P_{t}\zeta\sigma_{e}^{2}}{\sigma_{n}^{2}})$, $Z_{0}=\beta_{1}\sum_{n=1}^{N}|h_{IU,n}|^{2}$, and $Z_{1}=\beta_{0}\sum_{n=1}^{N}|H_{BI,m,n}|^{2}$. Under CLT approximation, $Z_{k},~k\in\{0,1\}$ follows Gaussian distribution, specifically $Z_{k}\sim\mathcal{N}(\mu_{z_{k}},\sigma^{2}_{z_{k}})$ with mean and variance as follows
$\mu_{z_{k}}$ $=\frac{Nc_{0}}{d_{k}^{\alpha_{k}}},$ (60a)
$\sigma^{2}_{z_{k}}$ $=\frac{Nc_{0}}{d_{k}^{2\alpha_{k}}}(\frac{(K_{k}^{2}+4K_{k}+2)}{(K_{k}+1)^{2}}-1),~\forall k.$ (60b)
@@ IV-B Active-IRS
In the case of CLT approximation, the OP of the Active-IRS system can be obtained as
@@
$R(x)=Q(\frac{\mu_{v}-\sqrt{\mu_{w}(x)}}{\sigma_{v}\sqrt{1-\rho^{2}}})-Q(\frac{\mu_{v}+\sqrt{\mu_{w}(x)}}{\sigma_{v}\sqrt{1-\rho^{2}}}),$ (62)
$\mu_{w}(x)=\frac{\gamma_{\text{th}}}{M\gamma_{0}x}(c_{1}\mu_{z_{0}}+c_{2}\mu_{z1}+c_{3}),$ (63)
$\rho=\frac{\mu_{vz}-\mu_{v}(c_{1}\mu_{z_{0}}+c_{2}\mu_{z1})}{\sigma_{v}\sqrt{c_{1}^{2}\sigma^{2}_{z_{0}}+c_{2}^{2}\sigma^{2}_{z1}}},$ (64)
$\mu_{vz}=\sqrt{\beta_{0}\beta_{1}(1-\zeta)}(c_{1}\beta_{1}A_{1}+c_{2}\beta_{0}A_{0}),$ (65)
$A_{p}=\frac{3\pi N}{8(1+K_{p})^{\frac{3}{2}}(1+K_{\bar{p}})^{\frac{1}{2}}}L_{\frac{3}{2}}(-K_{p})L_{\frac{1}{2}}(-K_{\bar{p}})$ (66)
$+\frac{\pi N(N-1)}{4(1+K_{0})^{\frac{1}{2}}(1+K_{1})^{\frac{1}{2}}}L_{\frac{1}{2}}(-K_{0})L_{\frac{1}{2}}(-K_{1}),\quad p\in\{0,1\}.$
@@
Let $Z=c_{1}Z_{0}+c_{2}Z_{1}$. Since two RVs $Z_{0}$ and $Z_{1}$ are independent, the RV of $Z$ follows Gaussian distribution, i.e., $Z\sim\mathcal{N}(\mu_{z},\sigma^{2}_{z})$, which $\mu_{Z}=c_{1}\mu_{z_{0}}+c_{2}\mu_{z_{1}}$ and $\sigma^{2}_{z}=c_{1}^{2}\sigma^{2}_{z_{0}}+c^{2}_{2}\sigma^{2}_{z_{1}}$. Now, we can rewrite SNR in (59) as
$\gamma_{\text{max}}=\frac{M\gamma_{0}\mathbb{G}U}{Z+c_{3}}.$ (67)
Therefore, the OP can be given as
$P_{\text{out}}=\int F_{\gamma_{\text{max}}|\mathbb{G}}(\gamma_{\text{th}})f_{\mathbb{G}}(\mathbb{G})d\mathbb{G},$ (68)
@@
$F_{\gamma_{max}|\mathbb{G}}(\gamma_{\text{th}})$ $=\text{Pr}(\frac{U}{Z+c_{3}}\leq\frac{\gamma_{\text{th}}}{M\gamma_{0}\mathbb{G}})$ (69a)
$=\int\text{Pr}(\frac{U}{z+c_{3}}\leq\frac{\gamma_{\text{th}}}{M\gamma_{0}\mathbb{G}}|Z=z)f_{Z}(z)dz$ (69b)
$=\int F_{U|Z}(\frac{\gamma_{th}(z+c_{3})}{M\gamma_{0}\mathbb{G}})f_{Z}(z)dz$ (69c)
$=\mathbb{E}_{Z}\{F_{U|Z}(w(\mathbb{G},z))\},$ (69d)
where $W(\mathbb{G},z)=\frac{\gamma_{th}(z+c_{3})}{M\gamma_{0}\mathbb{G}}$. The distribution of $W$ follows $\mathcal{N}(\mu_{w}(\mathbb{G}),\sigma_{w}^{2}(\mathbb{G}))$ where $\mu_{w}(\mathbb{G})=\frac{\gamma_{\text{th}}}{M\gamma_{0}\mathbb{G}}(c_{1}\mu_{z_{0}}+c_{2}\mu_{z_{1}}+c_{3})$ and $\sigma^{2}_{w}(\mathbb{G})=\frac{\gamma^{2}_{\text{th}}}{M^{2}\gamma^{2}_{0}\mathbb{G}^{2}}(c_{1}^{2}\sigma^{2}_{z_{0}}+c_{2}^{2}\sigma^{2}_{z_{1}})$. From (56), we have
$F_{U|Z}(w(\mathbb{G},z$ $))=$ (70)
$Q(\frac{\mu_{v|z}-\sqrt{w(\mathbb{G},z)}}{\sigma_{v|z}})-Q(\frac{\mu_{v|z}+\sqrt{w(\mathbb{G},z)}}{\sigma_{v|z}}).$
$V$ and $Z$ are jointly Gaussian with correlation coefficient $\rho$, the conditional mean $\mu_{v|z}$ and variance $\sigma^{2}_{v|z}$ are given by
$\mu_{v|z}$ $=\mu_{v}+\rho\frac{\sigma_{v}}{\sigma_{z}}(z-\mu_{z}),$ (71a)
$\sigma^{2}_{v|z}$ $=\sigma^{2}_{v}(1-\rho^{2}),$ (71b)
where $\rho=\frac{\text{Cov}(v,z)}{\sigma_{v}\sigma_{z}}$. Here, $\text{Cov}(v,z)=E\{vz\}-\mu_{v}\mu_{z}$ where $E\{vz\}$ is obtained as
$E\{vz\}=\sqrt{\beta_{0}\beta_{1}(1-\zeta)}(c_{1}\beta_{1}A_{1}+c_{2}\beta_{0}A_{0}).$ (72)
The terms $A_{p},p\in\{0,1\}$ are calculated as
$A_{p}=$ $\frac{3\pi N}{8(1+K_{p})^{\frac{3}{2}}(1+K_{\bar{p}})^{\frac{1}{2}}}L_{\frac{3}{2}}(-K_{p})L_{\frac{1}{2}}(-K_{\bar{p}})$ (73)
$+\frac{\pi N(N-1)}{4(1+K_{0})^{\frac{1}{2}}(1+K_{1})^{\frac{1}{2}}}L_{\frac{1}{2}}(-K_{0})L_{\frac{1}{2}}(-K_{1}).$
where $\bar{p}$ denotes the complement of $p$. Using (41), we approximate $F_{U|Z}(w(\mathbb{G},z))$ in (70) at point $\mu_{z}$ as follows
$F_{U|Z}(w(\mathbb{G},z))$ $\approx F_{U|Z}(w(G,\mu_{z}))$ (74)
$+F_{U|Z}^{(1)}(w(G,\mu_{z}))(z-\mu_{z})$
$+F_{U|Z}^{(2)}(w(G,\mu_{z}))(z-\mu_{z}))^{2}.$